a square peice of paper is folded in half. the perimeter of the resulting rectangle is 24 cm. find the area of the original square.
Answer by mananth(16946)

let the length of side of square be x
...
when you fold it into half
one side will be x and the other x/2
...
2(x+x/2)=24
x+x/2 = 24/2
x+x/2 = 12
(2x+x)/2 = 12
3x=24
x= 8 cm the original length of square.
...
m.ananth@hotmail.ca